If a rectangles perimiter is 40 and the length is 1 1/2 longer what are the length and width

Answer by Alan3354(69443)

If a rectangles perimiter is 40 and the length is 1 1/2 longer what are the length and width
---------------------
P = 2L + 2W = 40
L + W = 20
-----
L = W + 1.5
-----
W + W + 1.5 = 20
2W = 18.5
W = 9.25
L = 10.75